OpenStudy (anonymous):

-300/(-13) -300/-13=-300/13=23.07692308 was wrong The quotient of a neg and a pos value is neg, but the quotient of two neg and two posvalues is pos. Help! I do not understand what instructor means.

OpenStudy (diyadiya):

-ve / -ve =+( positive) -300/-13 = 300/13=23.07692308

OpenStudy (diyadiya):

For example \[\frac{-4}{2}= -2\]\[\frac{-4}{-2}= 2\]\[\frac{4}{2}=2\]

OpenStudy (anonymous):

That is the answer I had and she told me to recheck my work

OpenStudy (diyadiya):

-300/-13=-300/13 <--this is where you went wrong its not -300/13 its actually 300/13=23.07.......

OpenStudy (anonymous):

gotcha, thanks

OpenStudy (diyadiya):

Welcome